OpenStudy (anonymous):

Determine whether the graphs of the pair of lines are parallel. x+6=y y-x = -2 The slope of the line x+6=y is ? (or is it undefined)?

OpenStudy (anonymous):

that should be x+6=y and y-x = -2 I didn't mean for them to run together.

OpenStudy (anonymous):

y=x+6 and y=x-2 are your l ines Take the derivative (or look at the coefficient of x, slope, m) to find that y'=1 for both of them. They're parallel